A car moving at 50km/h skids 15m with locked brakes. How far will the car skid with locked brakes at 150km/h?
Answer by khwang(438)

Let the deceleration be a m/h^2.
Use solve v^2 = 2a S , 50^2 = 2a * 15 to find
a = 250/3 m/ h^2.

When v = 150, S = v^2/ 2a = 150^2 / (500/3) = (9 * 2500)* (3/500)
= 27* 5 = 135 meters.
Or use V1^2/S1 = V2^2/S2 and solve
50^2/15 = 150^2/S2
or 1/15 = 9/S2
to get S2 = 135 m.
Kenny